Circular No. 6899

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bu IN NGC 3368 Mirko Villi, Forlí, Italy, reports his discovery on CCD images taken on

May 9.9 UT of an apparent supernova (mag ∼ 13) located ∼ 60′′ north of the center of NGC 3368 = M96. Villi notes that a previous CCD image taken on Apr. 21.9 shows no object in this position, neither is anything visible on several atlases of galaxy images. S. Nakano, Sumoto, Japan, reports that M. Aoki (Toyama) has obtained a CCD image of SN 1998bu (V = 11.8) on May 10.618 with a 0.43-m reflector in bad weather conditions, from which the following precise position was measured: α = 10^{h}46^{m}46^{s}.03, δ = +11^{o}50′07′′.1 (equinox 2000.0), which is 4′′.3 east and 55′′.3 north of the center of NGC 3368. No star was visible on previous frames of this galaxy taken by Aoki. B. A. Skiff, Lowell Observatory, reports that he viewed the apparent supernova on May 10.33 UT at visual mag perhaps 12.0–12.5. D. Hanžl, Brno, Czech Republic, reports that he imaged SN 1998bu on May 10.835 at V = 12.37 with a 0.40-m reflector + CCD; he gave the offset as 49′′ north of the galaxy’s center.

GRB 980425 J. S. Bloom, California Institute of Technology (CIT); D. Frail, Na-

tional Radio Astronomy Observatory; and S. R. Kulkarni, CIT, report on behalf of the CIT γ-ray burst team: “Using the 2.54-m reflector at Las Campanas, S. Vogel and M. Regan obtained V, R, and I images of the field of GRB 980425 (IAUC 6884) near the presumed supernova discovered by Galama et al. (IAUC 6895, 6898). Using the secondary standards of Lidman et al. (IAUC 6895), we determine that the transient has continued to brighten, with the following derived magnitudes (± 0.05): May 8.311 UT, V = 13.87; 8.306, R = 13.84; 8.309, I = 13.98. The error comes predominately from the zero-point uncertainty. The contribution to the flux from a nearby star ∼ 4′′.9 to the northwest and the host galaxy is negligible (< 0.02 mag). A composite light curve from published fluxes can

be obtained at http://astro.caltech.edu/~jsb/ltcurve 980425.ps. It

is still unclear as to the connection between the presumed supernova and GRB 980425. In order to correlate the time of both, an accurate light curve must be obtained.”

1998 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green